# Tracing the bottom electroweak dipole operators at future lepton colliders

### v1

@@ IV Results
Figure 1: The preferred region from a global fit with $\delta g_{Lb}$, $\delta g_{Rb}$, $\kappa_{bA}$, $\kappa_{bZ}$ to the combination of the future Z-pole and $\sqrt{s}=240\text{ GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}=\left\{-3.80,-5.15,-0.102,-0.985\right\}\times 10^{-4}$.
In Fig. missing 1, we present the preferred regions in the parameter space of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$, obtained from a four-parameter global fit to the measurements at the Z-pole and 240 GeV. The results in the top panel are obtained by keeping only the linear dependence of the effective parameters in the observables, while for the bottom panel the full dependence is kept. For later convenience, we will refer the two scenarios by “linear fit” and “full fit”, respectively. In each case, the $\chi^{2}=1$ and $\chi^{2}=4$ contours are projected on the $(\kappa_{bA},\kappa_{bZ})$ plane (left panel) and the $(\delta g_{Lb},\delta g_{Rb})$ plane (right panel) by profiling over the other two parameters. The measurements at two different energies provide four independent constraints on the four parameters, and from the results we see that indeed the four parameters can be constrained simultaneously. On the other hand, we observe a visible difference between the results of the linear fit and the full fit, suggesting that the squared terms of the dipole contributions play an important role in the fit. The full fit results also exhibit a peculiar second best-fit point, where all the contributions from the four parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ cancel in the four observables, which takes the SM values. Such points are indeed possible since setting the four observables to SM values gives a set of non-linear equations that may have multiple solutions. Overall, the precision of the effective parameters is at the $10^{-4}$ level or better.
@@
$\delta g_{Lb}$$\pm 2.66$ 1 [-5.34, 1.87]
$\delta g_{Rb}$$\pm 3.17$ 0.919 1 [-7.94, 2.32]
$\kappa_{bZ}$$\pm 0.621$ 0.975 0.914 1 [-1.41, 0.398]
$\kappa_{bA}$$\pm 0.249$ 0.804 0.784 0.868 1 [-0.228, 0.228]
@@
$\delta g_{Lb}$$\pm 2.64$ 1 [-2.58, 1.50]
$\delta g_{Rb}$$\pm 3.15$ 0.918 1 [-3.30, 2.10]
$\kappa_{bZ}$$\pm 0.616$ 0.975 0.913 1 [-0.624, 0.304]
$\kappa_{bA}$$\pm 0.247$ 0.803 0.783 0.868 1 [-0.126, 0.168]
@@
V The LEP $A^{0,b}_{\rm FB}$ discrepancy
One explanation for the discrepancy between the LEP $A^{0,b}_{\rm FB}$ measurement and the SM prediction is to have nonzero $\delta g_{Lb}$ and $\delta g_{Rb}$ from new physics, which can be generated by e.g. introducing heavy vector-like quarks that mix with the bottom Choudhury:2001hs . It is interesting to see whether a sizable bottom dipole coupling could offer a viable alternative solution. To test this possibility, we focus on the current Z-pole measurements of $R_{b}$ and $A^{0,b}_{\rm FB}$ in Table 1.444Adding $A_{b}$ to the fit would slightly alter the central values and the bounds but would not qualitatively affect the results.
For the dipole couplings, we also focus on $\kappa_{bZ}$ and omit $\kappa_{bA}$, since $\kappa_{bZ}$ has the dominant contribution at the Z-pole. With two measurements $R_{b}$ and $A^{0,b}_{\rm FB}$, it is not possible to simultaneously constrain the three parameters $\delta g_{Lb},\,\delta g_{Rb}$ and $\kappa_{bZ}$. We thus consider two-parameter fits by setting either $\delta g_{Lb}$ or $\delta g_{Rb}$ to zero.
Figure 3: Left: Constraints from the LEP $R_{b}$ $A_{FB}$ measurements in the $(\kappa_{bZ},\delta g_{Lb})$ plane, setting $\delta g_{Rb}=0$ and keeping only the linear dependence of $\kappa_{bZ}$ and $\delta g_{Lb}$. Right: The same result but for $(\kappa_{bZ},\delta g_{Rb})$ with $\delta g_{Lb}$ set to zero. For each observable, the solid and dashed lines corresponds to the central value and the 1-$\sigma$ bounds, respectively. The best-fit point is at $\{\kappa_{bZ},\delta g_{Lb}\}=\{-0.00469,-0.0163\}$ on the left panel and $\{-0.000669,0.0163\}$ on the right one. The gray region corresponds to the combined $\chi^{2}$ of $R_{b}$ and $A^{0,b}_{\rm FB}$ being smaller than one. The black dot indicates the SM.
Figure 4: Same as Fig. missing 3 but with the full dependence of the two parameters in each panel.
Our results are shown in Fig. missing 3 which consider only the linear contributions, and Fig. missing 4 which include the full contributions (in particular the dipole-squared ones). With only linear contributions, there always exist a point in the parameter space that produces the LEP central values for both $R_{b}$ and $A^{0,b}_{\rm FB}$, as illustrated in Fig. missing 3. With dipole-squared contributions, the situation becomes more interesting. We note that the dipole amplitude-squared term alone has a zero asymmetry because the corresponding final states $b$ and $\bar{b}$ have the same helicity, and the distribution is thus symmetric under the exchange of the two particles. Therefore, having a non-zero dipole-squared contribution would make $A^{0,b}_{\rm FB}$ smaller than the SM value, in the direction favored by the LEP measurements (see Table 1). The results in Fig. missing 4 indeed verify this. For $\delta g_{Lb}=\delta g_{Rb}=0$ (corresponding to the gray horizontal line in both plots), two preferred regions of $\kappa_{bZ}$ with different signs can be found that are consistent with the LEP $A^{0,b}_{\rm FB}$ measurement. For a perfect fit with LEP $R_{b}$ and $A^{0,b}_{\rm FB}$ measurements, a nonzero $\delta g_{Lb}$ is also needed. (This could not be achieved with only $\delta g_{Rb}$ and $\kappa_{bZ}$, as shown on the right panel of Fig. missing 4). Nevertheless, we see that with $\kappa_{bZ}$ alone (setting $\delta g_{Lb}=\delta g_{Rb}=0$), one could already obtain a good agreement (within $\chi^{2}\lesssim 1$) with the LEP $R_{b}$ and $A^{0,b}_{\rm FB}$ measurements.
One potential challenge for the bottom dipole explanation is that the dipole operators are usually generated by loop contributions, while $\delta g_{Lb}$ and $\delta g_{Rb}$ could be generated by tree-level effects. Our results suggest that $|\kappa_{bZ}|\sim 0.001$ in order to generate the observed $A^{0,b}_{\rm FB}$. Assuming $c_{bW},c_{bB}\sim 1/16\pi^{2}$, Eq. (missing) 6 would then give $\Lambda\lesssim 10^{2}$ GeV, making SMEFT invalid. Nevertheless, the situation can be significantly improved in a strongly coupled scenario, in which case the bottom dipole remains to be a viable explanation for the $A^{0,b}_{\rm FB}$ discrepancy.
VI Summary
@@
In this paper, we performed a phenomenological analysis on the $e^{+}e^{-}\to b\bar{b}$ measurements at a future lepton collider to extract the precision reach on a set of dimension-6 bottom-quark operators. Focusing on CP-even effects, we found that the two bottom dipole operators, $\mathcal{O}_{bW}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)\sigma^{i}HW_{\mu\nu}^{i}$ and $\mathcal{O}_{bB}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)HB_{\mu\nu}$, could have sizable contributions to the $e^{+}e^{-}\to b\bar{b}$ process despite the suppression of the interference term with SM by the bottom mass. The combination of measurements at multiple center-of-mass energies is crucial for discriminating dipole operators from the ones that modify the SM $Zb\bar{b}$ couplings. The dipole quadratic contributions (from the square of the dipole amplitudes) are not suppressed by the bottom mass and their effects can be comparable with the linear ones. With measurements at only two energies (e.g. Z-pole and 240 GeV), the inclusion of the quadratic contributions leads to a non-SM second best-fit point in the parameter space due to cancellations among the linear and quadratic terms, which can be lifted by including the measurements at an additional energy (e.g. 360 GeV). Overall, we found that a future lepton collider, with measurements at both the Z-pole and higher energies, provides powerful probes on the bottom dipole operators, with a precision reach corresponding to a new physic scale in the multi-TeV range, assuming an order-one coupling. We also find that the long-standing discrepancy of the LEP $A^{0,b}_{\rm FB}$ measurement could also be caused by a non-zero dipole contribution. A future lepton collider will not only resolve the current $A^{0,b}_{\rm FB}$ discrepancy Gori:2015nqa , but also find out whether the discrepancy has contributions from bottom dipole operators should it stay.
@@
The importance of the dipole quadratic contributions in our analysis naturally raises concerns about the validity of the SMEFT framework, since they are at the $1/\Lambda^{4}$ level, formally the same as the linear contribution of dimension-8 operators. It should be noted that, for dimension-8 dipole operators, the interference term with the SM is also suppressed by the bottom mass. Treating the bottom mass as another small expansion parameter, one could indeed keep the quadratic contribution of dimension-6 dipole operators while omitting the linear dimension-8 ones. More generally, the robustness of the SMEFT analysis can be improved by either cutting the phase space where the quadratic contributions dominate Contino:2016jqw , or treating the quadratic contributions as theory error Alte:2017pme . Here we chose to simply present both the results with the quadratic contributions and the ones without, leaving a more delicate treatment of SMEFT validity to future studies.
@@
There are several directions in which our study can be extended. We have restricted the collider scenario to be a circular collider without beam polarizations. It is straightforward to apply the same analysis to linear colliders with beam polarizations. In addition to the possibility of directly measuring the asymmetry observable $A_{b}$, with different beam polarizations one could also effectively change the ratio between the $Z$ and photon diagrams, which receive different dipole contributions. As such, beam polarization could provide additional discrimination power in the fit. We have focused on the CP-even two-bottom operators in our study, and an obvious extension is to also consider the CP-odd effects. This introduces more parameters, and it could be challenging to find observables that separate the CP-odd effects from the CP-even ones.  [^5] This is unlike the top quark, for which one could exploit the top decay distribution. It is also important to implement our analysis in a more global framework, in particular with the top-quark operators, as already pointed out in Durieux:2019rbz ; Liu:2022vgo . We leave these important directions to future studies.
@@ Acknowledgments
We thank Yong Du, Christophe Grojean, Yiming Liu, Bin Yan for useful discussions and valuable comments on the manuscripts. This work is supported by the National Natural Science Foundation of China (NSFC) under grant No. 12035008, No. 12375091 and No. 12347171. X.Z. Tan also acknowledges the support from Helmholtz – OCPC (Office of China Postdoc Council) Postdoctoral Fellowship-Program.
@@ Appendix A Results in terms of the dim-6 Wilson coefficients
Here we present the results in terms of the dim-6 Wilson coefficients $(c_{Hq}^{(1)}+c_{Hq}^{(3)})$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$. The $1\sigma$ bounds and correlations are summarized in Table 4, while the preferred regions are shown in Fig. missing 5 (for Z-pole + 240 GeV) and Fig. missing 6 (for Z-pole + 240 GeV + 360 GeV). Note that the results here are the same as the ones in Section IV but only in a different basis.
@@
Figure 5: The preferred region from a global fit with $c_{Hq}^{(1)}+c_{Hq}^{(3)}$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$ to the combination of the Z-pole and $\sqrt{s}=240\text{ GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\ (c_{Hq}^{(1)}+c_{Hq}^{(3)}),\,c_{Hb},\,c_{bB},\,c_{bW}\right\}=\left\{-1.25,-1.70,-7.65,-11.33\right\}\times 10^{-2}$.
@@
Figure 6: Same as Fig. missing 5 but also includes the 360 GeV measurements. The second minimum in the full fit is lifted.

### v2

@@ IV Results
Figure 1: The preferred region from a global fit with $\delta g_{Lb}$, $\delta g_{Rb}$, $\kappa_{bA}$, $\kappa_{bZ}$ to the combination of the future Z-pole and $\sqrt{s}=240\,\text{GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=\chi^{2}_{\rm min}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}=\left\{-7.80,0.420,0.423,-1.22\right\}\times 10^{-4}$.
In Fig. missing 1, we present the preferred regions in the parameter space of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$, obtained from a four-parameter global fit to the measurements at the Z-pole and 240 GeV. The results in the top panel are obtained by keeping only the linear dependence of the effective parameters in the observables, while for the bottom panel the full dependence is kept. For later convenience, we will refer to the two scenarios by “linear fit” and “full fit”, respectively. In each case, the contours at the confidence level of 68% and 95% are projected on the $(\kappa_{bA},\kappa_{bZ})$ plane (left panel) and the $(\delta g_{Lb},\delta g_{Rb})$ plane (right panel) by profiling over the other two parameters. The measurements at two different energies provide four independent constraints on the four parameters, and from the results we see that indeed the four parameters can be constrained simultaneously. On the other hand, we observe a visible difference between the results of the linear fit and the full fit, suggesting that the squared terms of the dipole contributions play an important role in the fit. The full fit results also exhibit a peculiar second best-fit point, where all the contributions from the four parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ cancel in the four observables, which takes the SM values. Such points are indeed possible since setting the four observables to SM values gives a set of non-linear equations that may have multiple solutions. Overall, the precision of the effective parameters is at the $10^{-4}$ level or better.
@@
$\delta g_{Lb}$$\pm 2.65$ 1 [-8.90, 2.20]
$\delta g_{Rb}$$\pm 1.36$0.200 1 [-1.33, 1.68]
$\kappa_{bZ}$$\pm 0.375$ 0.975 -0.271 1 [-1.408, 0.304]
$\kappa_{bA}$$\pm 0.127$ 0.112 -0.088 0.201 1 [-0.118, 0.655]
@@
$\delta g_{Lb}$$\pm 2.63$ 1 [-2.88, 1.68]
$\delta g_{Rb}$$\pm 1.36$0.193 1 [-1.20, 1.20]
$\kappa_{bZ}$$\pm 0.372$ 0.975 -0.264 1 [-0.42, 0.225]
$\kappa_{bA}$$\pm 0.125$ 0.112 -0.088 0.202 1 [-0.078, 0.096]
@@
V Summary
@@
In this paper, we performed a phenomenological analysis on the $e^{+}e^{-}\to b\bar{b}$ measurements at a future lepton collider to extract the precision reach on a set of dimension-6 bottom-quark operators. Focusing on CP-even effects, we found that the two bottom dipole operators, $\mathcal{O}_{bW}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)\sigma^{i}HW_{\mu\nu}^{i}$ and $\mathcal{O}_{bB}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)HB_{\mu\nu}$, could have sizable contributions to the $e^{+}e^{-}\to b\bar{b}$ process despite the suppression of the interference term with SM by the bottom mass. The combination of measurements at multiple center-of-mass energies is crucial for discriminating dipole operators from the ones that modify the SM $Zb\bar{b}$ couplings. The dipole quadratic contributions (from the square of the dipole amplitudes) are not suppressed by the bottom mass, and their effects can be comparable with the linear ones. With measurements at only two energies (e.g. Z-pole and 240 GeV), the inclusion of the quadratic contributions leads to a non-SM second best-fit point in the parameter space due to cancellations among the linear and quadratic terms, which can be lifted by including the measurements at an additional energy (e.g. 360 GeV). Overall, we found that a future lepton collider, with measurements at both the Z-pole and higher energies, provides powerful probes on the bottom dipole operators, with a precision reach corresponding to a new physics scale in the multi-TeV range, assuming an order-one coupling.
@@
The importance of the dipole quadratic contributions in our analysis naturally raises concerns about the validity of the SMEFT framework, since they are at the $1/\Lambda^{4}$ level, formally the same as the linear contribution of dimension-8 operators. It should be noted that, for dimension-8 dipole operators, the interference term with the SM is also suppressed by the bottom mass. Treating the bottom mass as another small expansion parameter, one could indeed keep the quadratic contribution of dimension-6 dipole operators while omitting the linear dimension-8 ones. We also note that the Wilson coefficients that modify the SM $Zb\bar{b}$ couplings, $c^{(1)}_{Hq}+c^{(3)}_{Hq}$ and $c_{Hb}$, are much better constrained than the dipole coefficients, as shown in Table 4, and their square terms have little impact in our analysis. As such, we do not expect that the dimension-8 operators that modify the SM $Zb\bar{b}$ couplings are relevant to our analysis. It is indeed nontrivial to build a model that generates the bottom dipole interactions while keeping modifications to the SM $Zb\bar{b}$ couplings small. This is however not a major concern in our study, since we take a bottom-up approach assuming Wilson coefficients are free parameters subject to experimental constraints, as in many other SMEFT analyses. More generally, the robustness of the SMEFT analysis can be improved by either cutting the phase space where the quadratic contributions dominate Contino:2016jqw , or treating the quadratic contributions as the theory error Alte:2017pme . Here we chose to simply present both the results with the quadratic contributions and the ones without, leaving a more delicate treatment of SMEFT validity or the construction of viable UV models to future studies. We also note that the differences between the linear-fit results and the full-fit ones in Table 3 are relatively small (with at most $\sim 40\%$ relative differences) if all measurements up to 360 GeV are included. Therefore, we would expect the results to be still within the same order of magnitude, should a more strict treatment of EFT validity be applied (e.g. by treating the quadratic contributions as theory error).
@@
There are several directions in which our study can be extended. We restricted the collider scenario to a circular collider without beam polarizations. It is straightforward to apply the same analysis to linear colliders with beam polarizations. In addition to the possibility of directly measuring the asymmetry observable $A_{b}$, with different beam polarizations one could also effectively change the ratio between the $Z$ and photon diagrams, which receive different dipole contributions. As such, beam polarization could provide additional discrimination power in the fit. We have focused on the CP-even two-bottom operators in our study, and an obvious extension is to also consider the CP-odd effects. This introduces more parameters, and it could be challenging to find observables that separate the CP-odd effects from the CP-even ones.  [^5] This is unlike the top quark, for which one could exploit the top decay distribution. It is also important to implement our analysis in a more global framework, in particular with the top-quark operators, as already pointed out in Durieux:2019rbz ; Liu:2022vgo . We leave these important directions to future studies.
@@ Acknowledgments
The authors thank Yong Du, Christophe Grojean, Yiming Liu, Bin Yan for useful discussions and valuable comments on the manuscripts. X.Z. Tan also thanks Lukas Allwicher, Lisong Chen for discussions about quark running mass, and Guilherme Guedes especially for help on the UV model searching with the SOLD package although not being included ultimately in this work. This work is supported by the National Natural Science Foundation of China (NSFC) under grant No. 12035008, No. 12375091 and No. 12347171. X.Z. Tan also acknowledges the support from Helmholtz – OCPC (Office of China Postdoctoral Council) Postdoctoral Fellowship Program.
@@ Appendix A Results in terms of the dim-6 Wilson coefficients
Here we present the results in terms of the dim-6 Wilson coefficients $(c_{Hq}^{(1)}+c_{Hq}^{(3)})$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$. The $1\sigma$ bounds and correlations are summarized in Table 4, while the preferred regions are shown in Fig. missing 3 (for Z-pole + 240 GeV) and Fig. missing 4 (for Z-pole + 240 GeV + 360 GeV). Note that the results here are the same as the ones in Section IV but only in a different basis.
@@
Figure 3: The preferred region from a global fit with $c_{Hq}^{(1)}+c_{Hq}^{(3)}$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$ to the combination of the Z-pole and $\sqrt{s}=240\,\text{GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\ (c_{Hq}^{(1)}+c_{Hq}^{(3)}),\,c_{Hb},\,c_{bB},\,c_{bW}\right\}=\left\{-2.57,0.139,-2.70,-17.6\right\}\times 10^{-2}$. For convenience, $\Lambda$ is set to 1 TeV for all the coefficients here.
@@
Figure 4: Same as Fig. missing 3 but also includes the 360 GeV measurements. The second minimum in the full fit is lifted. $\Lambda$ is also set to 1 TeV here.
Appendix B Results by using running mass $m_{b}(m_{Z})=2.86\,\text{GeV}$
In Table 5 and 6, we present the fit results with combined Z-pole, 240 GeV and 360 GeV measurements, which uses the bottom mass value $m_{b}(m_{Z})=2.86\,\text{GeV}$ (in the $\overline{\rm MS}$ scheme with the renormalization scale $\mu=m_{Z}$). They should be compared with the (bottom panel of) Table 3 and 4, which use $m_{b}(m_{b})=4.18\,\text{GeV}$. For the linear fit, we expect the bounds for the dipole coefficients to be worsen by a factor of $m_{b}(m_{b})/m_{b}(m_{Z})\approx 1.5$, which is indeed the case for $c_{bB}$ and $c_{bW}$ in Table 6. For $\kappa_{bZ}$ and $\kappa_{bA}$ in Table 5, however, there appears to be little change in the linear fit result. This is merely due to the parameterization we use in Eq. (missing) 3, where $\kappa_{bZ}$ and $\kappa_{bA}$ are normalized by a factor of $m_{b}$. For the full fit, the change in the results are somewhat less predictive due to the nontrivial interplay between linear and quadratic terms. Nevertheless, the overall results for $c_{bB}$ and $c_{bW}$ are still worse for a smaller $m_{b}$ as expected.
Z-pole + 240 GeV + 360 GeV (with $m_{b}=2.86$ GeV)
Linear fit Correlation $\rho$ Full fit
$\delta g_{Lb}$$\pm 2.63$ 1 [-2.64, 1.52]
$\delta g_{Rb}$$\pm 1.34$0.188 1 [-1.20, 1.20]
$\kappa_{bZ}$$\pm 0.373$ 0.975 -0.260 1 [-0.390, 0.195]
$\kappa_{bA}$$\pm 0.125$ 0.115 -0.090 0.205 1 [-0.072, 0.186]
[…]
Z-pole + 240 GeV + 360 GeV (with $m_{b}=2.86$ GeV)
Linear fit Correlation $\rho$ Full fit
$(c_{Hq}^{(1)}+c_{Hq}^{(3)})$$\pm 0.869$ 1 [-0.871, 0.501]
$c_{Hb}$$\pm 0.444$0.188 1 [-0.396, 0.396]
$c_{bB}$$\pm 4.54$ 0.815 -0.246 1 [-3.12, 3.12]
$c_{bW}$$\pm 6.46$ 0.974 -0.249 0.778 1 [-8.16, 3.23]
[…]